OpenStudy (anonymous):

man i cannot figure this out. 2x^2-5x-3=0

OpenStudy (anonymous):

What exactly is your question?

OpenStudy (anonymous):

solve equation by factoring

OpenStudy (anonymous):

I think the answer is (x+3)(2x-1)

OpenStudy (anonymous):

Ok. We have\[2x^{2}-5x-3=0\]In order to solve this, let's remember that a quadratic is of form\[ax^{2}+bx+c\]In this case, we want to multiply a and c together:\[a \times c=2 \times -3 =-6\]We then want factors of ac that, when added together, equal b. The factors of -6 are: 6, -1 -6, 1 2, -3 -2, 3 We can see that -6+1=-5. So now we form this box:\[\left[\begin{matrix}2x^{2} & ? \\ ? & -3\end{matrix}\right]\]Notice that the first entry in the first row is ax^2, and that the second element of the second row is c. Now we're going to put in those factors of ac we found times x (it doesn't matter in which spot they go):\[\left[\begin{matrix}2x^{2} & -6x \\ 1x & -3\end{matrix}\right]\]Now let's find the greatest common factor of each row. In the first row, we could pull out 2x, so that's our first factor. In the second row, we can only pull out a 1, so that's our second factor. Now we're going to factor the columns. Firs the first column, the only common factor is x, so that's our first factor. In the second one though, both -6x and -3 are divisible by -3, so that is our second factor. Now, knowing those factors, we know that\[2x^{2}-5x-3=(x-3)(2x+1)=0\]Now that it's factored, can you finish the problem?

OpenStudy (anonymous):

is it -3/1,1/2 or -3,1/2

OpenStudy (anonymous):

It's the answers to these two equations:\[x-3=0\]\[2x+1=0\]

OpenStudy (anonymous):

oh that'll be 3, -1/2

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

thaks buddy youre a life saver

OpenStudy (anonymous):

thanks